OpenStudy (anonymous):

Please help me :(...The circle given by x^2+y^2-6y-12=0 can be written in standard form like this:x^2(y-k)^2=21...what is the value of k in this equation?

OpenStudy (shamim):

x^2+(y-3)^2=21

OpenStudy (anonymous):

could you explain more please

OpenStudy (shamim):

k=3

OpenStudy (anonymous):

you must complete the square: \[y^{2} + by = (y + b/2)^{2} - (b/2)^{2}\] \[y^2 - 6y = (y - 3)^{2} - (3)^{2}\]

OpenStudy (anonymous):

so it is 3 ?

OpenStudy (anonymous):

so the end equation is \[x^{2} + (y-3)^{2} - 9 - 12 = 0\] x^{2} + (y-3)^{2} = 21

OpenStudy (anonymous):

yes k = 3

OpenStudy (anonymous):

thanks guys,big help !

OpenStudy (anonymous):

glad i could help :) . remember and practice square completion. you'll use it later

OpenStudy (shamim):

i think the equation of a circle is \[x ^{2}+y ^{2}=a ^{2}\]